# Non-minimally coupled scalar field dark sector of the universe: in-depth (Einstein frame) case study

### v3

@@ 7 Bounce/cyclic evolution and its influence on the dynamical system
    Modified and quantum gravity formalisms [7, 90, 22].
@@ 8.4 Analytical initial conditions motivated by Planck 2018 and DESI DR2
    Slow-roll condition for Planck 2018 data (we only take into account the value of $\omega_{0}$):
@@
    Non-slow-roll CPL condition for Planck 2018 data (we take into account the value of $\omega_{0}$ and $\omega_{a}$):
@@
    Non-slow-roll CPL condition for DESI DR2+CMB+Pantheon+ (most geometric (background level) dataset) data (we take into account the value of $\omega_{0}$ and $\omega_{a}$):
@@
Additionally, if possible, we use the effective slope of the SF potential (see (8.34)) to determine possible values for the free parameters.
@@ 8.5 Model I: Axions/ALPs
The resulting evolution diagrams (Fig. 2-Fig. 4) demonstrate the following qualitative behavior:
@@ 8.6 Model II: Cyclic ekpyrotic
The resulting evolution diagrams (Fig. 5-Fig. 7) demonstrate the following qualitative behavior:
@@ 8.7 Model III: Exponential with constant
In this model, the resulting evolution diagrams (Fig. 8-Fig. 10) demonstrate the following qualitative behavior:
@@ 8.8 Model IV: Tracking quintessence
The resulting evolution diagrams (Fig. 11-Fig. 16) demonstrate the following qualitative behavior:
@@ 8.9 Model V: SFDM
The resulting evolution diagrams (Fig. 17-Fig. 19) demonstrate the following qualitative behavior:

### v4

# arXiv:2505.07456v4[gr-qc] 01 Mar 2026

@@ 7 Bounce/cyclic evolution and its influence on the dynamical system
    Modified and quantum gravity/cosmology formalisms [7, 103, 23, 70];
  4. (4)
    Higher-dimensional physics/cosmology (e.g., string theory and superstring theory [100, 101, 51, 52]) [64, 113, 49, 87, 33, 71].
@@ 8.4 Analytical initial conditions motivated by Planck 2018 and DESI DR2
    Slow-roll conditions for Planck 2018 data (we only take into account the value of $\omega_{0}$, see D for numerical plots):
@@
    Non-slow-roll CPL conditions for Planck 2018 data (we take into account the value of $\omega_{0}$ and $\omega_{a}$, see D for numerical plots):
@@
    Non-slow-roll CPL conditions for DESI DR2+CMB+Pantheon+ data (most geometric (background level) dataset) data (we take into account the value of $\omega_{0}$ and $\omega_{a}$):
@@
Additionally, if possible, we use the effective slope of the SF potential (8.34) to determine possible values for the free parameters. Comparing the numerical solutions obtained for the physical parameters may provide some clues about the ’inconsistencies’ in the results obtained in Planck 2018 and DESI DR2, which are related to cosmological tensions [42].
@@ 8.5 Model I: Axions/ALPs
The resulting evolution diagrams (Fig. 2, Fig. 8, Fig. 9) demonstrate the following qualitative behavior:
@@ 8.6 Model II: Cyclic ekpyrotic
The resulting evolution diagrams (Fig. 3, Fig. 10, Fig. 11) demonstrate the following qualitative behavior:
@@ 8.7 Model III: Exponential with constant
In this model, the resulting evolution diagrams (Fig. 4, Fig. 12, Fig. 13) demonstrate the following qualitative behavior:
@@ 8.8 Model IV: Tracking quintessence
The resulting evolution diagrams (Fig. 5, Fig. 6, Fig. 14 - Fig. 17) demonstrate the following qualitative behavior:
@@ 8.9 Model V: SFDM
The resulting evolution diagrams (Fig. 7, Fig. 18, Fig. 19) demonstrate the following qualitative behavior:
@@ Acknowledgements
The author also thanks the Reviewer for valuable comments, suggestions, and remarks on the manuscript.
@@ C.10 Extension to the full system
Appendix D Numerical evolution plots for the Planck 2018 initial conditions